### v2

@@ 6.2 Mechanisms
The results of this research indicate that sexual orientation anti-discrimination laws could significantly impact the labor market gaps for men in male same-sex partnerships. It is possible that these laws are binding and eliminate significant discrimination in the labor market. However, it is relatively easy for an employer to create a reason to fire an employee rather than firing someone specifically for being gay. It is also difficult to accurately determine someone’s sexual orientation by looking at them, unlike race and sex. Although it may be easier to determine someone’s sexual orientation conditional on them being in a same-sex relationship. Sexual orientation anti-discrimination laws may not be binding in a meaningful way that eliminates discrimination.
@@ 6.3 Threats to External Validity
One limitation of this research is that we are unable to view all LGB workers. The analysis is specifically conducted on individuals in same-sex partnerships and different-sex partnerships. The data does not allow for the identification of single LGB workers, whose outcomes may be significantly different than LGB workers in partnerships. Another pitfall to using partnerships to infer sexual orientation is that one could incorrectly infer someone’s sexual orientation from a partnership. Bisexual individuals represent a larger group than lesbian or gay individuals ((Gates, 2011)), and our specification erroneously classifies bisexual people in different-sex relationships as heterosexual. This misclassification biases estimates toward zero, as bisexual workers face significant wage penalties ((Mize, 2016)) and their inclusion in the reference group depresses the heterosexual average. Future work must disentangle these groups to avoid obscuring inequities.
Another potential concern in using partnerships to infer sexual orientation is that people in a same-sex partnership may be the group that is most at risk of discrimination. It seems likely that single LGB workers can more plausibly stay in the closet with their co-workers compared to their counterparts in a same-sex partnership. This analysis may be capturing the effect of anti-discrimination laws on the group that is most likely to be affected. These results may have a larger effect size than what one would find in examining the generalized LGB population.
These biases are for the broader LGB population and external validity. Notably, these potential biases do not affect internal validity. Translating the effect of anti-discrimination laws for people in same-sex partnerships to all other LGB workers is not immediately obvious. It is dependent on several factors that are unknowable in this analysis such as the wages and labor supply for single LGB workers and bisexual workers in different-sex partnerships and how discrimination affects LGB people for those in same-sex partnerships.
Furthermore, it is important to emphasize that this study cannot identify transgender individuals in the ACS, nor can it estimate the impact of anti-discrimination laws specifically protecting gender minorities. This restricts the ability to generalize findings to the entire LGBTQ+ community.
@@ 7 Conclusion
This analysis is the first quasi-experimental study examining the impact of both local and state anti-discrimination laws on labor supply and wage disparities between workers in same-sex and different-sex partnerships. Utilizing a novel city-level panel dataset capturing the granular detail of anti-discrimination laws, we find that these laws significantly reduce sexual orientation-based inequalities in labor market outcomes. Specifically, anti-discrimination laws decrease the labor force participation difference by 3.8 percentage points (78%) and employment difference by 3.7 percentage points (77%) for men in male same-sex partnerships, with insignificant effect for women in female same-sex partnerships. We also document important wage effects. Sexual orientation anti-discrimination laws lead to significant wage gains for men in same-sex partnerships, with their percentile rank in the hourly wage and annual earnings distribution improving on average by approximately 2.7 and 3 percentiles, respectively. Moreover, these laws increase the probability of ranking above the 25th percentile in hourly wages by 3.6 percentage points and above the 50th percentile by 3.4 percentage points, with similar improvements observed in the annual earnings distribution. By contrast, the wage distribution for women in same-sex partnerships remains largely unaffected by these laws.
We identify increased positive societal sentiments toward LGB Americans, reflected in growing support for same-sex marriage, as a potential mechanism through which these laws improve labor market outcomes. Furthermore, we explain differential responses to these laws between male and female same-sex households using Becker’s household specialization model and provide empirical support for this explanation.
Prior to the landmark Supreme Court decision in Bostock v. Clayton County (2020), many states permitted employers to fire employees based solely on sexual orientation. While Bostock extended federal protections under Title VII to LGBTQI+ workers, legislative momentum for anti-discrimination laws at the state level has slowed significantly since then. The Equality Act, reintroduced in the 118th Congress on June 21, 2023, seeks to expand federal civil rights laws to explicitly include sexual orientation and gender identity. As of March 2025, the bill is pending in the Senate, facing significant opposition from some senators, particularly concerning its implications for religious freedoms and women’s rights in certain contexts. At the state level, legislative momentum for anti-discrimination laws has slowed. Between 2009 and 2019, more states barred cities from protecting their residents from discrimination than those that expanded protections. Additionally, recent developments have seen states like Arkansas and Tennessee actively prohibiting localities from enacting their own anti-discrimination measures.
Our research provides critical and timely insights into the impact of anti-discrimination laws on labor market outcomes for individuals in same-sex partnerships. By offering a comprehensive evaluation of these laws’ effects, our findings meaningfully inform ongoing policy discussions and judicial decisions, including deliberations by the Supreme Court of the United States, about extending employment protections to the LGB community.
Appendix A Additional Robustness Checks
A.1 Event Studies
Figures A1, A2, A3, A4, A5 present the event study regressions corresponding to the robustness checks discussed in Sections 5.2, 5.3, 5.4, 5.5, 5.6. In each figure, we trace the evolution of labor supply impacts on the external margin relative to the enactment of anti-discrimination laws. Across all specifications, the pre-treatment coefficients are statistically insignificant, and the post-treatment dynamics closely mirror those observed in our main analyses. Specifically, estimates for women in metropolitan areas are smaller on average, unweighted models and those estimated before the legalization of same-sex marriage typically produce slightly lower coefficients, and both the post-2008 sample and non-movers yield somewhat larger effects.
A.2 Same-Sex Marriage Fixed Effects
@@
The results in Table A1 reinforce the robustness of our primary findings for men. Although the extensive margin effects for men in same-sex partnerships are slightly lower (0.0267 for labor force participation and 0.0270 for employment) than in our primary specification, they remain statistically significant, with generally smaller intensive margin estimates. For women, the point estimates on the extensive margin are quite similar in Table A1 and Table 6. However, the employment effect for women turns negative and is marginally significant (–0.0212, p $<$ 0.10), which contrasts with the mostly null effects observed elsewhere, aside from the negative effects noted in metropolitan areas. This conflicting evidence suggests that further research is needed to better understand the impacts on women. Our hypothesis remains that high child-care costs in metropolitan areas may be an important contributing factor.
A.3 Occupation Fixed Effects
LGB workers tend to sort to different occupations and industries based on sex and sexual orientation [Black et al., 2007]. They show that gay men tend to sort to more female-dominated fields and lesbian women sort to more male-dominated fields relative to their heterosexual counterparts. It is typical to not consider those factors because discrimination affects occupation and industry choice, which would make occupation and industry choice endogenous.
Although controlling for occupation can pose selection concerns, because individuals not in the labor force typically lack occupation codes, we nonetheless examine within-occupation effects of anti-discrimination laws on wages and labor supply. Specifically, we employ regressions with four-digit occupation fixed effects to assess how wage and labor supply disparities shift following the enactment of these policies. [^18] In order to estimate within-occupation effects while still capturing individuals who are not in the labor force, we include a separate “no occupation” category for those without valid occupational codes. This approach allows us to retain non-employed respondents in the analysis, rather than excluding them outright due to missing occupation information. We acknowledge that categorizing all non-employed individuals into a single group can mask heterogeneity in reasons for not working (e.g., unemployment, home production, or retirement). Nonetheless, by creating a distinct “no occupation” category, we preserve a measure of occupational affiliation for those who are employed while ensuring that non-employed individuals remain part of the sample. One potential mechanism is that workers in male same-sex partnerships may transition into more in-demand occupations, thereby mitigating relative labor supply and wage gaps. However, as shown in Tables A2 and A3, even after accounting for occupation, men in same-sex partnerships exhibit significant changes in both labor supply and wages. These findings suggest that the impact of anti-discrimination laws is not merely a by-product of occupational sorting. Nonetheless, caution is warranted when interpreting these results, given the inherent limitations of conditioning on occupation in a broader labor force context.
Appendix Tables

### v3

@@ 6.3 Mechanisms
The results suggest that sexual-orientation anti-discrimination laws can meaningfully reduce labor-market gaps for men in male same-sex couples. One interpretation is that these laws are binding and help limit discrimination. At the same time, enforcing such protections is fundamentally different from enforcing protections based on race or sex, since sexual orientation is not always observable. These laws may become more binding when individuals are in publicly identifiable same-sex relationships, but less so otherwise. [^16] This idea is consistent with research on the saliency of protected class membership, such as armour2018disability, which shows that anti-discrimination protections are more effective when group membership is more readily perceived. Thus, while the evidence points to important effects for male same-sex couples, the extent to which anti-discrimination laws eliminate broader forms of sexual-orientation discrimination remains uncertain.
@@ 6.4 Threats to External Validity
A primary limitation of this study is that we cannot observe the full LGB population. The analysis compares individuals in same-sex partnerships to those in different-sex partnerships, but the ACS does not identify single LGB workers, whose labor market outcomes may differ systematically from partnered LGB individuals. Partnership status is also an imperfect proxy for sexual orientation: bisexual individuals in different-sex relationships are classified as heterosexual, and some gay or lesbian workers may similarly appear in the control group. Because lesbian women tend to earn more than heterosexual women, gay men tend to earn less than heterosexual men, and bisexual workers earn less than both (mize2016sexual), such misclassification does not shift group means in a uniform direction. While the bias in level pay gaps is generally toward zero, the sign of the bias in the estimated ENDA coefficients is theoretically ambiguous—especially if bisexual workers respond differently to anti-discrimination protections than gay or lesbian workers. Misclassification that places gay or lesbian individuals into the control group reduces the contrast between groups and thus produces the usual attenuation mechanism. These issues highlight the need for future work that can distinguish gay, lesbian, and bisexual workers more precisely to avoid masking heterogeneity across subpopulations.
A related concern is that individuals in same-sex partnerships may constitute the subset of LGB workers most vulnerable to discrimination. Single LGB workers may be more able to conceal their sexual orientation in the workplace, whereas partnered individuals face greater visibility. The effects documented here may therefore reflect treatment effects for the group most exposed to discrimination, potentially overstating impacts for the broader LGB population.
Crucially, these limitations pertain to external validity rather than internal validity. Misclassification of sexual orientation does not affect the within-county, within-sex comparisons that identify $\beta_{1g}$, but it does restrict the generalizability of the results to groups not directly observed in the data—such as single LGB workers or bisexual individuals in different-sex partnerships. How ENDA protections translate to these populations depends on their baseline outcomes, patterns of disclosure, and exposure to discrimination, none of which can be observed in the ACS.
Finally, the ACS does not allow us to identify transgender or nonbinary individuals, nor does it permit analysis of anti-discrimination laws that specifically protect gender minorities. This further limits generalizability to the broader LGBTQ+ community.
@@ 7 Conclusion
The policy environment has evolved considerably in recent years. Prior to Bostock v. Clayton County (2020), many states allowed employees to be fired solely based on sexual orientation. Although Bostock extended federal Title VII protections to LGBTQI+ workers, new anti-discrimination statutes have stagnated at both the state and federal levels. The Equality Act, reintroduced in the 118th Congress on June 21, 2023, remains pending in the Senate, with critics voicing concerns related to religious liberties and women’s rights. Meanwhile, more states have chosen to preempt or restrict local anti-discrimination ordinances than those that have adopted new protections, reflecting persistent polarization.
Against this backdrop, our study offers quasi-experimental evidence on how local and state-level anti-discrimination laws shape labor market outcomes for individuals in same-sex couples. Exploiting a novel city-level panel dataset that captures the granular evolution of these laws, we document significant reductions in sexual orientation-based labor market disparities. In particular, for men in male same-sex couples, our estimates indicate a 3.8 percentage-point (78%) decrease in the labor force participation gap and a 3.7 percentage-point (77%) decline in the employment gap relative to different-sex counterparts, whereas we find no significant changes for women in female same-sex couples.
Beyond labor supply, our results reveal meaningful wage improvements for men in same-sex couples. Their average percentile rank in hourly wages and annual earnings increases by about 2.6 and 2.9 points, respectively, with a 3.7 percentage-point increase in the likelihood of surpassing the 25th wage percentile and a 3.3 percentage-point boost for exceeding the 50th percentile. These outcomes underscore the capacity of anti-discrimination laws to mitigate wage gaps rooted in sexual orientation, with no comparable impact observed for women in female same-sex couples.
We identify increased positive societal sentiments toward LGB Americans, reflected in growing support for same-sex marriage, as a potential mechanism by which these laws improve labor market outcomes. Taken together, our results suggest that local and state-level anti-discrimination measures can significantly mitigate labor market disadvantages faced by men in same-sex couples. By illustrating how these laws reduce labor supply and wage gaps, we offer timely insight into ongoing policy debates and judicial deliberations about extending employment protections to sexual minorities, paving the way for future research on the broader implications of such measures.
Appendix A County Identification from PUMAs
The ACS reports Public Use Microdata Areas (PUMAs) rather than counties. We assign respondents to counties using the IPUMS PUMA–county crosswalk. In practice, metro PUMAs map nearly one-to-one to counties: 475 counties in our sample are uniquely identified. Rural PUMAs aggregate multiple counties within a state; in these cases, we follow ACS public-use geography and treat each rural PUMA as a single county-group. This yields 48 county-groups, for a total of 523 county (or county-group) units. All ENDA codes are constructed at the county (or county-group) $\times$ year level using this mapping.
Appendix B Using Only State Laws
Local ordinances provide meaningful identifying variation. Among the 332 counties in our sample that ever receive ENDA-type protection, 133 (about 40 percent) adopt a county-level ordinance before their state enacts a statewide ENDA. A state-only measure would classify all residents in these county–years as untreated, which affects roughly 30 percent of post-treatment ACS person-year observations. Ignoring county ordinances therefore induces non-classical measurement error in the treatment variable and mechanically attenuates the estimated effects.
Table B.1 illustrates this attenuation by removing local-law variation. In Panel A, which replaces county fixed effects with state fixed effects, point estimates shrink relative to our baseline (Table 4), and, importantly, the male–female differences on the extensive margin become statistically insignificant. Panel B, which keeps the original fixed-effects structure but forces ENDA to vary only at the state level, shows a similar pattern: some effects shrink towards zero, and the male–female gap is less precisely estimated.
Table B.1: Effect of Anti-Discrimination Laws: Extensive & Intensive Margin of Labor Supply (State Laws)
(1) (2) (3) (4) (5) (6)
Extensive Margin Intensive Margin
Panel A: State-Fixed Effects
ENDA $\times$ SSP: Female 0.0040 -0.0089 1.5019∗∗ -0.3407 1.0750∗∗∗ -1.4526∗∗∗
(0.0136) (0.0189) (0.5864) (0.7646) (0.2932) (0.3630)
ENDA $\times$ SSP: Male 0.0270∗∗ 0.0243 0.2947 -0.1546 -0.2827 -1.0244∗∗∗
(0.0125) (0.0125) (0.3015) (0.5411) (0.3488) (0.3585)
$p$-value: M - F 0.2831 0.2062 0.0793 0.8527 0.0053 0.3204
Panel B: County-Fixed Effects
ENDA $\times$ SSP: Female -0.0067 -0.0158 1.5380∗∗∗ -0.3971 1.2360∗∗∗ -1.3461∗∗∗
(0.0106) (0.0140) (0.5051) (0.6383) (0.3126) (0.3535)
ENDA $\times$ SSP: Male 0.0326∗∗ 0.0321 0.6748∗∗ 0.2175 -0.2064 -0.7743∗∗
(0.0153) (0.0169) (0.3292) (0.5922) (0.5124) (0.3419)
$p$-value: M - F 0.0294 0.0754 0.1967 0.5485 0.0117 0.1645
Employed Only X X
Observations 7,781,358 7,781,358 7,781,358 7,781,358 5,765,580 5,765,580
Mean (F) 0.818 0.782 34.42 39.51 41.25 47.27
Mean (M) 0.825 0.790 35.53 39.96 42.41 47.69
  • Notes: Data is derived from the American Community Survey (ACS) conducted annually between 2005 and 2019. We compare individuals in same-sex partnerships with those in different-sex partnerships within four years of ENDA enactment. For this table, we abstract from local laws and use only state-level ENDA variation. The regressions are estimated jointly, allowing the treatment effect to differ for men and women through interaction terms. Columns (1–2) focus on the extensive margin, examining changes in labor force participation and employment. Columns (3–6) explore the intensive margin using variation in weeks worked per year and usual weekly hours; Columns (5–6) restrict the sample to employed individuals. The “Mean (F)” and “Mean (M)” rows report pre-treatment averages of the outcome variables (including never-treated counties) for women and men in same-sex couples, respectively. Reported $p$-values test whether the ENDA effect differs significantly between men and women. Panel A replaces county fixed effects with state fixed effects in Equation (1), while Panel B reports the original specification. Standard errors are clustered at the state-level. $p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
These results confirm that omitting local ordinances discards meaningful within-state variation and leads to substantial attenuation of the estimated ENDA effects, particularly on the extensive margin where the gender differences weaken considerably.
Appendix C Additional Robustness Checks
C.1 Placebo Tests and Alternative Estimators
We next conduct a series of placebo tests and apply alternative estimators to assess the impact of anti‑discrimination laws on labor market outcomes among both same‑sex and different‑sex couples. Specifically, we utilize four distinct methods to assess the impact of anti‑discrimination laws on labor market outcomes among same‑sex and different‑sex couples: TWFE, callaway2021difference, sun2021estimating, and borusyak2021revisiting. As a placebo test, we apply the DiD framework to different‑sex couples, who are not the target of these laws, and thus we expect no significant changes in their labor market outcomes. Panel (a) of Tables C.1 and C.2 report the DiD estimates for different‑sex couples, while Panels (a) and (c) of Figures C.1 and C.2 display the corresponding event study plots.
Next, Panel (b) of Tables C.1 and C.2 presents the impact of anti‑discrimination laws on same‑sex couples, while Panels (b) and (d) of Figures C.1 and C.2 display the corresponding event studies. In these analyses, we compare labor market outcomes for same‑sex couples in regions that have enacted the laws with those in regions that have not. The results reveal that male same‑sex couples in areas with anti‑discrimination laws experience significant improvements in labor market outcomes, particularly through increased labor force participation and higher employment rates. For female same‑sex couples, however, we observe no significant effects on the extensive margin; instead, the positive and significant impact is confined to the intensive margin, as evidenced by an increase in average weekly hours worked. Hence, our alternative estimators yield similar point estimates and comparable confidence intervals, reinforcing the robustness of our findings.
((a)) Labor Force Participation (Female Different-Sex Couples)
((b)) Labor Force Participation (Female Same-Sex Couples)
((c)) Labor Force Participation (Male Different-Sex Couples)
((d)) Labor Force Participation (Male Same-Sex Couples)
Figure C.1: Impacts on Labor Force Participation by Partnership Type
Notes: Event study plots on the effects of ENDAs on people in same-sex couples and different-sex couples broken down by sex following the county-level and state-level anti-discrimination laws. Coefficients with 95% confidence intervals. Standard errors are clustered at the county level. Reference year is -4.
((a)) Employment (Female Different-Sex Couples)
((b)) Employment (Female Same-Sex Couples)
((c)) Employment (Male Different-Sex Couples)
((d)) Employment (Male Same-Sex Couples)
Figure C.2: Impacts on Employment by Partnership Type
Notes: Event study plots on the effects of ENDAs on people in same-sex couples and different-sex couples broken down by sex following the county-level and state-level anti-discrimination laws. Coefficients with 95% confidence intervals. Standard errors are clustered at the county level. Reference year is -4.
Table C.1: Effect of Anti-Discrimination Laws on Men in Same-Sex and Different-Sex Partnerships
(1) (2) (3) (4) (5) (6)
Extensive Margin Intensive Margin
ENDA $\times$ SSP Estimates Labor Force Employed Weekly Hours Weeks Worked Weekly Hours Weeks Worked
Panel A: Men in Different-Sex Partnerships
TWFE 0.0004 -0.0002 0.0629 0.0642 -0.0022 -0.0233
(0.0015) (0.0018) (0.0834) (0.0809) (0.0602) (0.0448)
callaway2021difference 0.0008 -0.0012 0.1038 0.0596 0.0797 0.0016
(0.0013) (0.0017) (0.0961) (0.0817) (0.0725) (0.0595)
sun2021estimating 0.0008 -0.0005 0.1160 0.0827 0.0349 -0.0239
(0.0015) (0.0019) (0.0852) (0.0911) (0.0570) (0.0496)
borusyak2021revisiting 0.0023 0.0006 0.1898∗∗ 0.1345 0.0554 -0.0415
(0.0014) (0.0020) (0.0879) (0.0958) (0.0552) (0.0465)
Observations 3,763,442 3,763,442 3,763,442 3,763,442 3,046,376 3,046,376
Employed Only X X
Panel B: Men in Same-Sex Partnerships
TWFE 0.0284∗∗ 0.0263 0.8832 1.1614 -0.1199 -0.2060
(0.0137) (0.0144) (0.5618) (0.6476) (0.4351) (0.3773)
callaway2021difference 0.0463∗∗ 0.0428∗∗ 1.2572∗∗ 1.3140 0.1736 -0.1863
(0.0184) (0.0213) (0.6368) (0.8613) (0.5461) (0.4090)
sun2021estimating 0.0337∗∗ 0.0332∗∗ 0.9484 1.3165∗∗ -0.1495 -0.1319
(0.0145) (0.0152) (0.5357) (0.6657) (0.3888) (0.3290)
borusyak2021revisiting 0.0420∗∗∗ 0.0414∗∗∗ 1.3754∗∗∗ 1.7464∗∗∗ -0.0756 -0.1261
(0.0129) (0.0137) (0.5001) (0.6246) (0.3827) (0.2779)
Observations 35,194 35,194 35,194 35,194 27,504 27,504
Employed Only X X
  • Notes: Data is derived from the American Community Survey (ACS) conducted annually between 2005 and 2019. We compare men in counties with ENDAs with those without within four years of ENDA enactment. The regressions are estimated separately by partnership: Panel A for men in different-sex partnerships, while Panel B for men in same-sex partnerships. The coefficients indicate the impact of anti-discrimination laws. The analysis is divided into two parts: Columns (1-2) focus on the extensive margin, examining changes in labor force participation and employment status as a result of the laws. Columns (3-6), on the other hand, explore the intensive margin by leveraging variations in the number of weeks worked per year and average weekly working hours. Standard errors are clustered at the county level. Results are shown using six different estimators: Two-Way Fixed Effects (TWFE), callaway2021difference, sun2021estimating, and borusyak2021revisiting. $p<0.10$, ∗∗ $p<0.05$, $^{{}^{***}}$ $p<0.01$.
Table C.2: Effect of Anti-Discrimination Laws on Women in Same-Sex and Different-Sex Partnerships
(1) (2) (3) (4) (5) (6)
Extensive Margin Intensive Margin
ENDA $\times$ SSP Estimates Labor Force Employed Weekly Hours Weeks Worked Weekly Hours Weeks Worked
Panel A: Women in Different-Sex Partnerships
TWFE 0.0010 0.0008 0.0093 0.0045 -0.0882 -0.1083∗∗
(0.0019) (0.0017) (0.0781) (0.0756) (0.0566) (0.0485)
callaway2021difference 0.0017 0.0020 -0.0127 -0.0105 -0.0693 -0.0602
(0.0019) (0.0023) (0.1021) (0.1016) (0.0733) (0.0783)
sun2021estimating 0.0010 0.0010 -0.0152 -0.0159 -0.0732 -0.0705
(0.0017) (0.0018) (0.0710) (0.0830) (0.0618) (0.0602)
borusyak2021revisiting 0.0015 0.0013 0.0057 -0.0130 -0.0637 -0.1021
(0.0016) (0.0016) (0.0733) (0.0845) (0.0620) (0.0641)
Observations 3,942,881 3,942,881 3,942,881 3,942,881 2,660,191 2,660,191
Employed Only X X
Panel B: Women in Same-Sex Partnerships
TWFE 0.0099 0.0090 1.3686∗∗∗ 0.4999 0.5733 -0.4873
(0.0092) (0.0105) (0.4386) (0.5327) (0.3579) (0.3743)
callaway2021difference -0.0042 -0.0088 0.9758 -0.6803 0.8353 -1.2359∗∗
(0.0120) (0.0125) (0.7503) (0.6246) (0.4648) (0.4891)
sun2021estimating 0.0022 -0.0009 1.1645∗∗ -0.0565 0.5538 -0.8987∗∗
(0.0103) (0.0109) (0.4920) (0.5785) (0.3148) (0.3529)
borusyak2021revisiting 0.0000 -0.0013 1.0359∗∗ -0.2617 0.5542 -0.9571∗∗∗
(0.0111) (0.0129) (0.4939) (0.5924) (0.2977) (0.3228)
Observations 40,071 40,071 40,071 40,071 31,707 31,707
Employed Only X X
  • Notes: Data is derived from the American Community Survey (ACS) conducted annually between 2005 and 2019. We compare women in counties with ENDAs with those without within four years of ENDA enactment. The regressions are estimated separately by partnership: Panel A for women in different-sex partnerships, while Panel B for women in same-sex partnerships. The coefficients indicate the impact of anti-discrimination laws. The analysis is divided into two parts: Columns (1-2) focus on the extensive margin, examining changes in labor force participation and employment status as a result of the laws. Columns (3-6), on the other hand, explore the intensive margin by leveraging variations in the number of weeks worked per year and average weekly working hours. Standard errors are clustered at the county level. Results are shown using six different estimators: Two-Way Fixed Effects (TWFE), callaway2021difference, sun2021estimating, and borusyak2021revisiting. $p<0.10$, ∗∗ $p<0.05$, $^{{}^{***}}$ $p<0.01$.
C.2 Same-Sex Marriage Fixed Effects
@@
The results in Table C.3 reinforce the robustness of our primary findings for men. Although the extensive-margin effects for men in same-sex couples are slightly lower (0.0267 for labor force participation and 0.0270 for employment) than in our primary specification, they remain statistically significant, with generally smaller intensive-margin estimates. For women, the point estimates on the extensive margin are quite similar in Table C.3 and Table 4, but the employment effect turns negative and is marginally significant (0.0212-0.0212, $p<0.10$), which contrasts with the mostly null extensive-margin effects observed elsewhere, aside from the negative responses noted in metropolitan areas. Moreover, the joint model indicates that the ENDA effects on the extensive margin are significantly larger for men than for women, as reflected in the $p$-values reported in Columns (1) and (2). This conflicting evidence for women suggests that further research is needed to better understand the mechanisms at play; our hypothesis remains that high child-care costs in metropolitan areas may be an important contributing factor.